OpenStudy (anonymous):

If 2^ (16x^2 - 8x-3) =1, what does x equal? a *1/4 and 3/4 b 3/4 only c 1/4 only d 1/4 and 3/4

OpenStudy (anonymous):

2^0 = 1 so 16^2 - 8x - 3 = 0 (4x-3)(4x+1) = 0 x = 3/4 or -1/4 2^(16(3/4)^2 -8(3/4) -3) = 1?? 2^(9 - 6 - 3) = 1?? 2^(0) = 1 ?? 1 = 1

OpenStudy (anonymous):

@LegoMyEgo its a equation all together

OpenStudy (anonymous):

What do you need help on?

OpenStudy (anonymous):

nvm i didnt understand at first thank you